OpenStudy (anonymous):

what is the square root of 2^3?

OpenStudy (anonymous):

@nyamamoto It is Biology Portion . Post your question in proper section.

OpenStudy (abhisar):

\(\sf \sqrt{2^3}=\huge 2^{\frac{3}{2}}=?\)

OpenStudy (anonymous):

well to answer your question nyamamoto it is 2.8284271247461900976033774484194

OpenStudy (abhisar):

\(\huge \sqrt{8}=2.82\)

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

this is biology dude...